OpenStudy (anonymous):

Find a function that describes the sequence. Use y for the term in the sequence and n for its place in the list. Then use the function to find the next three terms in the sequence. 7.5, 8.5, 9.5, 10.5, . . . a. y = n + 6.5; 11.5, 12.5, 13.5 c. y = 6.5n +1; 33.5, 40, 46.5 b. y = 7.5n; 37.5, 45, 52.5 d. y = 8.5n –1; 41.5, 50, 58.5

OpenStudy (anonymous):

Hmmph, typo. sec.

OpenStudy (anonymous):

ehm. By our sequence we know that first item is 7.5. So we can basically say: \[ n=1 \implies y=7.5 \] let's define \[ y=n+k\] Now let's say. In case that n = 1: \[ n=1 \implies y=1+k=7.5 \\ 1+k=7.5 \\ k=6.5 \\ y=n+6.5 \] Now to calculations. We have first 4 items, let's find fifth, sixth and seventh \[ n=5 \implies y=5+6.5=11.5 \\ n=6 \implies y=6+6.5=12.5 \\ n=7 \implies y=7+6.5=13.5 \]